Please help me solve the rest of this math problem, I am lost. I've done most of it.
Rite-Cut riding lawnmowers obey the demand equation p= -1/20x+1,030. The cost of producing x lawnmowers is given by the function C(x)= 150x+3,000.
a. Express the revenue R as a function of x. Simplify, do not factor answer.
The answer I got was 1,030x-1/20x^2
b. Express the profit P as a function of x. Simplify, do not factor answer
The answer I got was -1/20x^2+880x-3000
c. Find the value of x that maximizes profit. What is the maximum profit?
d. What price should be charged in order to maximize profit?

Answer by ewatrrr(24785)

c) P' = (-1/10)x + 880 = 0, x = 8800 (maz x-value)
P(8800) = $3,869,000
d) p = (-1/20)8800 + 1030 = $590, price to charge in order to maximize profit*